OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. x+y=3 2x+2y=10 I think the answer is no solutions

OpenStudy (anonymous):

you are right , these are parallel lines

OpenStudy (anonymous):

I think you are correct.

OpenStudy (anonymous):

y=-x+3 , y=-x+5

OpenStudy (anonymous):

sub y = 3-x in second equn. 2x + 2(3-x) = 10 2x -2x + 6 = 10 6 - 10 parallel lines there are infinite solutions

OpenStudy (anonymous):

jimmy is right, of course :)